OpenStudy (anonymous):

a sample of gas collected at 750.mm Hg occupies 250.mL at constant temperature what pressure does the gas exert if the volume increses to 300.mL

OpenStudy (whpalmer4):

You don't accept messages, but you send me messages asking for help? What's wrong with this picture?

OpenStudy (whpalmer4):

\[PV = nrT\]or for constant temperature, \[PV = k\]Figure out \(k\) from the pressure and volume at the time of collection, and use it to find the new pressure at the new volume.

OpenStudy (anonymous):

I didn't know i wasn't accepting messages

OpenStudy (anonymous):

this is confusing

OpenStudy (whpalmer4):

You know that the pressure and volume at the time of collection were 750 and 250 respectively. \(k = 750*250\) Now use that value of \(k\) along with your new volume of \(300\) to find the new pressure: \[750*250 = p*300\]Solve for \(p\). Units are the same as they were

OpenStudy (anonymous):

how do i solve for p

OpenStudy (whpalmer4):

same way you would solve \[a*b = c*p\]for \(p\)

OpenStudy (whpalmer4):

divide both sides by the thing being multiplied by \(p\)

OpenStudy (anonymous):

p=625

OpenStudy (anonymous):

is that right?

OpenStudy (whpalmer4):

if 750*250 = 625*300, then that is correct

OpenStudy (anonymous):

no it = 187500 but you have to divide by 300